1. If disposable income is 4,000, consumption is 3,500, government spending is 1,000, and taxes minus transfers are 800, national saving is equal to:

a. 300. b. 500. c. 700. d. 1,000.

2. Assume that equilibrium GDP (Y) is 5,000. Consumption (C) is given by the equation C= 500 + 0.6Y. Investment (I) is given by the equation I= 2,000 – 100r, where r is the real interest rate in percent. No government exists. In this case, the equilibrium real interest rate is:

a. 2 percent. b. 5 percent. c. 10 percent. d. 20 percent.
Answer #1

1. National savings = Public savings + Private savings

Private savings= Disposable income - Consumption

So here Private savings = 4000-3500=500

Public savings = (taxes net of transfers) - Government spending

So here public savings = 800-1000=-200

Thus national savings= 500+(-200)=300

Thus correct answer is option (a)

2. In equilibrium : Y=C+I

Given Y=5000 and C=500+0.6Y. Thus here C=500+0.6*5000=3500. Also given I=2000-100r

Substituting the above in the equilibrium condition we get : 5000=3500+2000-100r. Solving for r we get the equilibrium interest rate r=500/100=5

Thus correct answer is option (b)